Find the general solution of the DE
y² lny = x py+p².
Does the equation has any singular solution? If yes, obtain it.
1
Expert's answer
2019-08-20T09:32:36-0400

Let's rewrite the equation in form

"\ln y=x\frac{p}{y}+\frac{p^2}{y^2}"

Denote "Y=\ln y" hence "P=\frac{dY}{dx}=\frac{p}{y}"

Terefore

"Y=xP+P^2"

which is in the Clairaut's form


The solution can be obtained by just replacing P by constant c.

"Y=cx+c^2"

or

"\ln y=cx+c^2"

"y=\exp(cx+c^2)"


To find a singular solution, let's solve the equation

"(P^2)'+x=0"

"2P=-x"

"P=-\frac{x}{{2}}"

Therefore

"Y=-\frac{x^2}{2}+(-\frac{x}{2})^2=-\frac{1}{4}x^2"

will be the singular solution of this equation